OpenStudy (anonymous):

How do you solvev|2x+6|=8

OpenStudy (anonymous):

2x+6=8 -6 -6 2x=2 divide that and your answer is 1

OpenStudy (anonymous):

^-^ you should like become my tutor you help me understand it better

OpenStudy (anonymous):

5 dollars an hour xD aha jk but no worries ill try my best to give it to you step by step (: